# How do Laws Produce the Future?

### v1

@@ 8 Conclusion
The theories of gravity and electromagnetism that one encounters in introductory physics courses appear to be theories of time evolution that fit well with a dynamic production account of laws of nature. This paper has sought precision as to how the laws of these theories should be formulated so that they can be understood as producing the future from the past, guiding the ontology in its evolution. That exercise is important for understanding our options regarding the foundations of these theories and for assessing the merits of the dynamic production account of laws. The regular and relativistic time evolution paradigms have been put forward to clarify how laws produce the future and we have seen that Newtonian gravity and classical electromagnetism can be fit into these paradigms. To further develop the dynamic production account, the next step would be to build on the preliminary remarks in section 7 to see how more advanced physical theories might be fit into these paradigms.

### v2

@@ 8 Conclusion
The theories of gravity and electromagnetism that one encounters in introductory physics courses appear to be theories of time evolution that fit well with a dynamic production account of laws of nature. This paper has sought precision as to how the laws of these theories should be formulated so that they can be understood as producing the future from the past, guiding the ontology in its evolution. That exercise is important for understanding our options regarding the foundations of these theories and for assessing the merits of the dynamic production account of laws. The regular and relativistic time evolution paradigms have been put forward to clarify how laws produce the future and we have seen that Newtonian gravity and classical electromagnetism can be fit into these paradigms. To further develop the dynamic production account, one could build on the preliminary remarks in section 7 to see how more advanced physical theories might be fit into these paradigms.